OpenStudy (anonymous):

Solve y over negative 2+ 2 = 8

OpenStudy (anonymous):

\[\frac{y}{-4}=8?\]

OpenStudy (anonymous):

\[\frac{y}{-2} +2=8\] is it like this?????

OpenStudy (anonymous):

@adrian1998

OpenStudy (anonymous):

for \[-\frac{ y }{ 2 } + 2 = 8\] (If -2...) y - 4 = -16 (multiply both sides by -2) solve ^ that equation.